# Ensemble Debiasing Across Class and Sample Levels for Fairer Prompting Accuracy

### v1

@@ Abstract
Language models are strong few-shot learners and achieve good overall accuracy in text classification tasks, masking the fact that their results suffer from great class accuracy imbalance. We believe that the pursuit of overall accuracy should not come from enriching the strong classes, but from raising up the weak ones. To address the imbalance, we propose a post-hoc nonlinear integer programming based debiasing method that ensembles weight correction and membership correction to enable flexible rectifications of class probabilities at both class and sample levels, enhancing the performance of LLMs directly from their outputs. Evaluations with Llama-2-13B on seven text classification benchmarks show that our approach achieves state-of-the-art overall accuracy gains with balanced class accuracies. The resulted probability correction scheme demonstrates that sample-level corrections are necessary to elevate weak classes. In addition, due to effectively correcting weak classes, our method also brings significant performance gains to Llama-2-70B, especially on a biomedical domain task, demonstrating its effectiveness across both small and large model variants.
@@ 1 Introduction
Language models are good few-shot learners, but prompting can result in inevitable class accuracy imbalance, where a language model performs significantly better on certain classes compared to others, leading to biased performances across different classes. This issue is particularly relevant in tasks like classification, or any scenario where the model must distinguish between multiple categories. For example, on DBpedia ontology classification [1], the overall accuracy for 14 classes can reach 88% when prompting Llama-2-13B [2], but class Nature’s accuracy is 28%. Broadly, for open-ended generation tasks, when treating each token in the vocabulary as a distinct class, generating the most likely next token can be alternatively viewed as classifying over the entire vocabulary or soft labels [3, 4], so they are also susceptible to class accuracy imbalance. Therefore, this issue could be a prevalent problem affecting end users of many language model applications, and classification is the most basic case manifesting such imbalances.
@@
In text classifications, the prevalence of class accuracy imbalance is sometimes hidden by a single or a dominant evaluation metric, such as overall accuracy. Especially for relatively large models with over 1B parameters, the fact that they excel on benchmarks which emphasize overall performances may hide a poor performance in some classes. To enable a focused evaluation of whether language model performances are balanced across classes, the COBias metric was proposed [5] to dedicate evaluating pairwise class accuracy differences.
The issue is not particularly easy to solve from the root, because the causes span across training and prompting. First, language models often rely on priors learned during training. Model architecture built on the attention mechanism may be over-sensitive to certain input features, introducing various biases. If certain classes or concepts were more frequent or emphasized in the training data, the model may disproportionately favor them [6]. Secondly, the way prompts are phrased can unintentionally bias the model towards certain classes [7, 8, 9]. For example, ambiguous or leading prompts may cause the model to default to more common or “safer” predictions. In few-shot or zero-shot prompting, imbalanced class accuracy may be exacerbated because of underrepresented classes in the prompt demonstrative examples. Still, it helps to a limited extent by making all classes well-represented in the prompt demonstration [5]. Moreover, some classes may inherently be harder to distinguish due to semantic overlap or spurious correlations. For example, in entailment recognition, Entailment and Neutrality might be harder to differentiate when hypothesis sentences contain spurious triggers like negation words, leading to false Contradictions [10]. Considering the quantity of parameters in a language model, it is costly and inflexible to retrain the whole model to mitigate inherent data problems. Prompt engineering works case-by-case and does not adequately address the issue in most cases. Although a language model may provide the optimal response by generating self-correcting Chain-of-Thought outputs without relying on prompt engineering or finetuning [11], the emerging abilities to elicit correct reasoning paths still require thoughtfully designed reward signals. Lastly, particular types of biases such as gender bias or truthfulness can be traced back to internal representations and intervened at inference time [12, 13, 14], but layer-/attention-head-level interventions may not always be available to address the broader prediction bias in classification.
@@
To cost-effectively prevent the model from overly relying on learned patterns or knowledge, and to incorporate information present in downstream tasks, post-hoc optimization methods are introduced to address imbalanced accuracies in output classes. As such, post-hoc probability correction complements training and prompting strategies to improve model performances without altering the decoding procedures. It involves optimizing ICL output probabilities in inference time with a relatively small mathematical model, no matter what information and patterns a model has internalized from its training data and what human priors have been encoded in a prompt.
@@
Related Work. Prior post-hoc correction work has investigated re-weighting ICL output probabilities through combinatorial optimization addressing COBias[5, 15] or calibration addressing biased predictions [16, 6]. The former uses labeled instances for optimization while the latter does not. Both lines of work advance overall classification accuracy with low costs and the flexibility of no language model parameter updates. The performance difference is that the combinatorial optimization works achieve better balanced class performances while improving overall accuracy. However, their drawbacks lie in the single consideration at either class level [5] or sample level [15]. In [5], any sample gets re-weighted by the same set of per-class correction weights, making it hard to capture sample-level variances, which is particularly needed for weak classes. In [15], corrections are made at sample level by assigning per-sample per-class correction weights that are computed by membership functions, however, capturing too much nuance may instead over-corrects strong classes.
@@
To overcome the insufficiency in previous works and push forward higher prediction accuracy, we propose a meticulous post-hoc correction framework DCS (Debiasing at Class and Sample levels) that is an organic integration of both broad and fine-grained corrections. For classes that need to discriminate sample-level variances, we use fuzzy rules to perform targeted corrections; for those classes whose output probabilities do not vary across samples, we use simple class weights. For the rest of the paper, we first introduce the mathematical model for DCS ensemble debiasing (Section 2) and the overall framework (Section 3), followed by quantitative evaluations and in-depth analyses (Section 4). In Section 5, we conclude with prospective applications and open problems.
@@
2 Methodology
We introduce a robust nonlinear integer programming based framework to ensemble debiasing at both class and sample levels, which is targeted specifically for balancing class accuracy imbalances while improving overall accuracy over existing methods. To achieve ensembling, a Heaviside step function is designed to seamlessly unify two unique methods of correcting ICL outputs, bringing together the advantages of weight coefficient based correction and fuzzy rule based correction. This unified framework captures the intricacies of ICL output class probabilities with greater precision and clarity, making it possible to reach solutions that elevate weak classes while maintaining the strong ones. List of notations are summarized in Appendix A.
2.1 The Mapping Function for Probability Correction
@@
$\mu_{k}(p_{mi})=\left\{\begin{aligned} &0,&&p_{mi}\leq a_{k}\\ &\frac{p_{mi}-a_{k}}{b_{k}-a_{k}},&&a_{k}\leq p_{mi}\leq b_{k}\\ &\frac{c_{k}-p_{mi}}{c_{k}-b_{k}},&&b_{k}\leq p_{mi}\leq c_{k}\\ &0,&&\text{otherwise}\end{aligned}\right.$ (2)
@@
$p^{\prime}_{mi}=\mu_{\xi_{i}}(p_{mi})\cdot H(D_{F}-\xi_{i})+\omega_{\xi_{i}}(p_{mi})\cdot H(\xi_{i}-D_{F}-1)$ (4)
@@ 2.2 Mathematical Model
s.t. $\boldsymbol{\xi}=[\xi_{1},\dots,\xi_{N}],\xi_{i}\in\{1,\dots,D_{F}+D_{W}\}$ (5)
During optimization, at the beginning of each simulated annealing iteration, candidate selection $\xi_{i}$’s are generated and the correction functions $f_{i}$’s correct each sample in the optimization set, yielding corrected predictions $\{\hat{y}_{m}\}_{m=1}^{M}$. Each of the individual term in $Z$ is essentially an evaluation score computed using the corrected predictions and ground-truth labels $\{y_{m}\}_{m=1}^{M}$. If the new $Z$ is smaller than the current $Z$, the candidate selection is accepted and the best selection is updated; otherwise, the algorithm decides whether to accept the new selection based on the acceptance probability. Iterations repeat until a stopping criterion is met. The optimization process will be elaborated in Section 3.
@@ 3 Simulated Annealing for Solving DCS
Due to the non-differentiable nature of the math model, e.g., the use of argmax operators, it is challenging to solve directly using gradient based optimizers or standard OR (Operations Research) optimization solvers. To address this non-differentiable optimization problem, we must turn to heuristics. In this section, we employ the simulated annealing (SA) algorithm to solve the math model. The SA algorithm was proposed by [17, 18]. The forerunners successfully introduced the concept of annealing into combinatorial optimization, inspired by the annealing process of solid materials. It starts at an initial temperature and cools down at a specified rate. As the temperature gradually decreases, the algorithm leverages probabilistic jumps to explore the solution space randomly, searching for the global optimum of the objective function, so it can probabilistically escape local optima and eventually converge to the global optimum [19]. The optimization of DCS model is a search problem in multidimensional space, where each solution makes a potential probability correction scheme. Next, we uncover the overall framework that apply SA to solve our model.
@@ 3.1 Overall Framework
The overall solution framework for the DCS problem is depicted in Fig. 1. The combinatorial optimization relies on a labeled set of these initial class probabilities and their ground-truth labels. The labeled set can be simply taken from the full or a subset of the classification task’s training set. After ICL output class probabilities are obtained, the optimization process can start. It starts with an initial solution. An initial temperature $T$ with an annealing schedule indicating the cooling rate and the stopping criteria is defined. In each iteration, a neighborhood solution is generated by perturbing the current solution. The algorithm calculates the objective function value for the new solution (Step Evaluate Z). If the new solution is better, accept it; if the new solution is worse, accept it with a probability determined by the Metropolis criterion. Then the temperature is reduced according to the cooling rate. The iteration repeats until either a minimum temperature is reached or a maximum number of iterations is reached, and we obtain the optimal correction functions. In actual inference, the selected functions directly rectify a test instance’s ICL output class probabilities and result in corrected predictions.
@@ 3.2 Initial Solution
The initial solution can also be randomly generated, where $N$ random numbers are generated from $\mathbb{D}=\{1,\dots,D_{F}+D_{W}\}$ as the initial values for the selection variables $\xi_{j}$’s. For SA, the initial solution (how close it is to the optimal solution) does not affect the quality of the final solution, but it does influence the annealing process [17], and we empirically start with unchanged probabilities.
@@ 3.3 Neighborhood Solution
It is necessary to generate a new solution in the neighborhood of the current solution. If the current solution is $\xi_{j}=k_{j},\text{ for class }j=1,\dots,N$, we can randomly select a class, for example, $j^{\prime}$ whose current solution is $\xi_{j^{\prime}}=k_{j^{\prime}}$, and perturb it by sampling $k_{j^{\prime}}^{\text{New}}$ from the set $\mathbb{D}\setminus k_{j^{\prime}}$ as the new value for the variable $\xi_{j^{\prime}}$. That is, $\xi_{j^{\prime}}=k_{j^{\prime}}^{\text{New}}$. In this way, we obtain a neighboring solution: $\boldsymbol{\xi}^{\text{New}}=[\xi_{1}=k_{1},\dots,\xi_{j^{\prime}}=k_{j^{\prime}}^{\text{New}},\dots,\xi_{N}=k_{N}]$. The transition from the current solution at a given temperature to the new solution can be understood as a Markov chain [17, 18], and the transition probability is determined by the Metropolis criterion $P_{\text{accept}}=e^{-\Delta Z/T}$. If the new solution is better ($\Delta Z<0$), accept the new solution; if the new solution is worse ($\Delta Z\geq 0$) and a uniform random number $r\sim U(0,1)$ is smaller than $P_{\text{accept}}$, namely $r<P_{\text{accept}}$, the new solution is also accepted.
@@ 4 Experiments
We assess the capabilities of DCS in reducing class accuracy differences and enhancing overall accuracy with three primary ICL settings: 1-shot, 5-shot, and $N$-shot. Demonstrations in the prompt are randomly selected from the optimization set for 1-shot and 5-shot cases, while the $N$-shot demonstration cascades examples from each class, to minimize under-representations of any single class. The proposed method is compared against two related SoTA methods, DNIP and FuRud. We prompt LLMs with an A100 to obtain initial ICL outputs, following DNIP’s hyperparameters, then we run simulated annealing on CPUs. Evaluations are performed on seven multi-class text classification tasks across general and biomedical domains, including AGNews (AGN, 4-class;[20]), DBpedia (DBP, 14-class;[1]), SST-5 (SST, 5-class;[21]), TREC (6-class;[22, 23]), RTE (binary;[24]), DDI (5-class;[25]), and PubMedQA (PMQA, 3-class;[26]). See Appendix B for more details.
@@ 4.1 Main Results
The proposed method achieves SoTA overall accuracy improvements on all tasks except for RTE across all ICL settings. As for RTE, due to its simplicity as a binary classification task, all three methods obtain similar accuracy and COBias improvements. Meanwhile, it achieves SoTA COBias performance on general domain tasks. In general, its COBias is on par with the top methods, with average COBias over all seven tasks falling between that of DNIP and FuRud. On AGNews, DBpeida, and TREC, the proposed method achieves both more overall accuracy improvements and more COBias reduction than DNIP and FuRud across all ICL settings, demonstrating the superior performance gains achieved by ensembling class-level and sample-level debiasing. Weak classes are elevated. We show in Section 4.2 that DCS powers up the low-accuracy classes. We observe that weaker classes are often corrected by rules, while stronger classes are most likely corrected by weights, further highlighting the need of an ensemble method.
@@
Figure 2 plots the time used to optimize correction functions (left axis) and the search space size (right axis) for each task. For fair comparisons, we fix the weight scale for weight correction to 30 alongside 19 triangular membership functions, making a total of 49 possible correction functions and $49\cdot N$ possible assignments across all variables. DCS is computationally efficient where all tasks finish optimization within 150 outer loops, ranging from a maximum of 96.6 seconds to a minimum of 5.4 seconds. Notably, SST-5, DDI, and AGNews have a relatively longer time due to having a larger optimization set (Appendix B). In general, although larger optimization sets and search spaces need more time to converge, the annealing steps needed for a task can be further tuned to achieve cost-effective gains on downstream applications.

### v2

@@ Abstract
Language models are strong few-shot learners and achieve good overall accuracy in text classification tasks, masking the fact that their results suffer from great class accuracy imbalance. We believe that the pursuit of overall accuracy should not come from enriching the strong classes, but from raising up the weak ones. To address the imbalance, we propose a Heaviside step function based ensemble debiasing method, which enables flexible rectifications of in-context learned class probabilities at both class and sample levels. Evaluations with Llama-2-13B on seven text classification benchmarks show that our approach achieves state-of-the-art overall accuracy gains with balanced class accuracies. More importantly, we perform analyses on the resulted probability correction scheme, showing that sample-level corrections are necessary to elevate weak classes. Due to effectively correcting weak classes, our method also brings significant performance gains to a larger model variant, Llama-2-70B, especially on a biomedical domain task, further demonstrating the necessity of ensemble debiasing at both levels.
@@ 1 Introduction
Language models are good few-shot learners, but prompting can result in inevitable class accuracy imbalance. A language model can perform significantly better on certain classes compared to others, leading to biased performances across different classes (illustrated in Figure 1). This issue is particularly relevant in tasks like classification, or any scenario where the model must distinguish between multiple categories. For example, on DBpedia ontology classification [1], the overall accuracy for 14 classes can reach 88% when prompting Llama-2-13B [2], but class Nature’s accuracy is 28%. Broadly, for open-ended generation tasks, when treating each token in the vocabulary as a distinct class, generating the most likely next token can be alternatively viewed as classifying over the entire vocabulary or soft labels [3, 4], also susceptible to class accuracy imbalance. Therefore, accuracy imbalance is prevalent and affecting users of language model applications, where classification is the most basic case manifesting such imbalances.
Figure 1: The accuracy imbalance and what we can achieve with debiasing
@@
In text classifications, the prevalence of class accuracy imbalance is sometimes hidden by a single or a dominant evaluation metric, such as overall accuracy. Especially for relatively large models, the fact that they excel on benchmarks which emphasize overall performances may hide a poor performance in some classes. To enable a focused evaluation of whether language model performances are balanced across classes, the COBias metric was proposed [5] to dedicate evaluating pairwise class accuracy differences.
The class accuracy imbalance issue is not particularly easy to solve from the root. Its causes come from within the model architecture. In details, language models rely on priors learned during training to make predictions. However, built on self-attention, the models are prone to the copying mechanism of the induction heads [6, 7]. These heads predict repeated sequences, which can over-capture surface-level patterns in the data, leading to biased priors, and then biased predictions. Therefore, if certain classes or concepts were more frequently emphasized in the training data, the model may disproportionately favor them [8]. Although a systemic change at model level or selecting high-quality data may help, it can be inefficient.
Besides what is rooted in the model and data, the way prompts are phrased can further bias the model towards certain classes [9, 10, 11]. In few-shot prompting, imbalanced class accuracy can be exacerbated because of underrepresented classes in the demonstrations of a prompt. Nevertheless, it only helps to a limited extent by making all classes well-represented in the demonstrations [5]. In addition, Chain-of-Thought style outputs may facilitate self-correcting responses without relying on prompt engineering [12], but the outcome is guided and restricted by reward signals. Lastly, although certain types of biases, such as gender bias or truthfulness, can be traced back to Transformer layers and intervened at inference time [13, 14, 15], interventions can be inefficient to mitigate the accuracy bias, especially when it could be better treated without altering the model’s internal structure.
@@
Therefore, to effectively and efficiently reduce class accuracy imbalances, post-hoc optimization methods are introduced. As such, post-hoc probability correction complements training and prompting strategies to prevent the model from overly relying on learned patterns, and to incorporate necessary information in downstream tasks. This type of correction involves correcting In-context learned (ICL) output probabilities in inference time by combinatorially optimizing a relatively small mathematical model. These direct rectifications are useful, because they can directly correct outputs, no matter what biases are encoded in the model prior, or what prompts are used.
@@
Related Work. Prior post-hoc correction work has investigated re-weighting ICL output probabilities through combinatorial optimization addressing COBias[5, 16] or calibration addressing biased predictions [17, 8]. The former uses labeled instances for optimization while the latter does not. Both lines of work advance overall classification accuracy with low costs and the flexibility of no language model parameter updates. The performance difference is that the combinatorial optimization works achieve better balanced class performances while improving overall accuracy. We follow the combinatorial optimization perspective for our ensemble debiasing.
For a background on combinatorial optimization, it provides a mathematical framework for maximizing or minimizing an objective function of decision variables, subject to constraints on some or all of the variables, involving inequalities, equalities, and integrality restrictions [18]. This framework essentially transforms a rich variety of tasks into discrete optimization models, which is robust and remarkably versatile. It solves real-world problems such as planning, scheduling, etc. in operations research, and has diverse applications in NLP, from structured prediction such as named entity recognition and dependency parsing, to reading comprehension, then to grammatical error correction, and to align LLM outputs with human preferences [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. The core use of the framework is to translate a problem into a mathematical statement of one of the forms of linear mixed-integer programming problem (MIP), integer programming problem (IP), nonlinear integer programming problem (NIP), or other types [32, 33, 34, 30]. Solving combinatorial optimization models is not an easy task. Among which, NIP is highly challenging, because the NIP model can contain discontinuous functions and is inherently non-differentiable. Lin and You [5, 16] proposed to address the search difficulty with metaheuristics, which we follow (more discussions in Appendix A). Back to the combinatorial optimization based debiasing methods, their drawbacks lie in the single consideration at either class level [5] or sample level [16]. In [5], any sample gets re-weighted by the same set of per-class correction weights, making it hard to capture sample-level variances, which is particularly needed for weak classes. In [16], corrections are made at sample level by assigning per-sample per-class correction weights that are computed by membership functions, however, capturing too much nuance may instead over-corrects strong classes.
@@
To overcome the insufficiency in previous works, we propose a meticulous post-hoc correction framework DCS (Debiasing at Class and Sample levels). DCS is an organic integration of broad and fine-grained corrections across class and sample levels, pushing forward fairer prompting accuracy for LLMs. DCS automatically detects which classes need corrections that account for sample-level differences and which do not. In the following, we first introduce the Heaviside step function based ensembling strategy for DCS (Section 2) and the overall framework (Section 3), followed by quantitative evaluations and in-depth analyses (Section 4). In Section 5, we conclude with prospective applications and open problems.
@@
2 Using Heaviside Step Function to Ensemble Class and Sample Level Debiasing
We introduce a robust ensemble debiasing framework across class and sample levels, which is targeted specifically for balancing class accuracy while improving overall accuracy over individual methods. To achieve ensembling, a Heaviside step function seamlessly unifies two unique ICL output correction methods, bringing together the advantages of weight coefficient and fuzzy rule based corrections. This unified framework captures the intricacies of ICL output class probabilities with greater precision, reaching solutions that elevate weak classes while maintaining the strong ones. List of notations are presented in Appendix B.
2.1 The Mapping Function for ICL Probability Correction
@@
$\mu_{k}(p_{mi})=\left\{\begin{aligned} &0,&&p_{mi}\leq a_{k}\\ &\frac{p_{mi}-a_{k}}{b_{k}-a_{k}},&&a_{k}<p_{mi}\leq b_{k}\\ &\frac{c_{k}-p_{mi}}{c_{k}-b_{k}},&&b_{k}<p_{mi}\leq c_{k}\\ &0,&&\text{otherwise}\end{aligned}\right.$ (2)
@@
$p^{\prime}_{mi}=f_{i}(p_{mi})=\mu_{\xi_{i}}(p_{mi})\cdot H(D_{F}-\xi_{i})+\omega_{\xi_{i}}(p_{mi})\cdot H(\xi_{i}-D_{F}-1)$ (4)
@@ 2.2 Mathematical Model
s.t. $\boldsymbol{\xi}=(\xi_{1},\dots,\xi_{N}),\xi_{i}\in\{1,\dots,D_{F}+D_{W}\}$ (5)
where $1-Z^{\text{Acc}},Z^{\text{COBias}},Z^{\text{PMI}}$ respectively corresponds to the three terms of Eq. 5 in [5]. During optimization, at the beginning of each simulated annealing iteration, candidate selection $\xi_{i}$’s are generated and the correction functions $f_{i}$’s correct each sample in the optimization set, yielding corrected predictions $\{\hat{y}_{m}\}_{m=1}^{M}$. Each of the individual term in $Z$ is essentially an evaluation score computed using the corrected predictions and ground-truth labels $\{y_{m}\}_{m=1}^{M}$. If the new $Z$ is smaller than the current $Z$, the candidate selection is accepted and the best selection is updated; otherwise, the algorithm decides whether to accept the new selection based on the acceptance probability. Iterations repeat until a stopping criterion is met. The optimization process will be elaborated in Section 3.
@@ 3 Simulated Annealing for Solving DCS
Due to the non-differentiable nature of the math model, we turn to heuristics and employ the simulated annealing (SA) algorithm to solve the DCS model. The SA algorithm was proposed by [35, 36]. The forerunners successfully introduced the concept of annealing into combinatorial optimization, inspired by the annealing process of solid materials. It starts at an initial temperature and cools down at a specified rate. As the temperature gradually decreases, the algorithm leverages probabilistic jumps to explore the solution space randomly, searching for the global optimum of the objective function, so it can probabilistically escape local optima and eventually converge to the global optimum [37]. The optimization of DCS model is a search problem in multidimensional space, where each solution makes a potential probability correction scheme. Next, we uncover the overall framework that apply SA to solve our model.
@@ 3.1 Overall Framework
The overall solution framework for DCS is depicted in Fig. 2. The combinatorial optimization relies on a labeled set of these initial class probabilities and their ground-truth labels. The labeled set can be simply taken from the full or a subset of the classification task’s training set. After ICL output class probabilities are obtained, the optimization process starts with an initial solution. The iteration repeats until either a minimum temperature is reached or a maximum number of iterations is reached, and we obtain the optimal correction functions. In actual inference, the selected functions directly rectify a test instance’s ICL output class probabilities and result in corrected predictions.
@@ 3.3 Neighborhood Solution
It is necessary to generate a new solution in the neighborhood of the current solution. If the current solution is $\xi_{j}=k_{j},\text{ for class }j=1,\dots,N$, we can randomly select a class, for example, $j^{\prime}$ whose current solution is $\xi_{j^{\prime}}=k_{j^{\prime}}$, and perturb it by sampling $k_{j^{\prime}}^{\text{New}}$ from the set $\mathbb{D}\setminus k_{j^{\prime}}$ as the new value for the variable $\xi_{j^{\prime}}$. That is, $\xi_{j^{\prime}}=k_{j^{\prime}}^{\text{New}}$. In this way, we obtain a neighboring solution: $\boldsymbol{\xi}^{\text{New}}=(\xi_{1}=k_{1},\dots,\xi_{j^{\prime}}=k_{j^{\prime}}^{\text{New}},\dots,\xi_{N}=k_{N})$. The transition from the current solution at a given temperature to the new solution is a Markov chain [35, 36], and the transition probability is determined by the Metropolis criterion $P_{\text{accept}}=e^{-\Delta Z/T}$. If the new solution is better ($\Delta Z<0$), accept the new solution; if the new solution is worse ($\Delta Z\geq 0$) and a uniform random number $r\sim U(0,1)$ is smaller than $P_{\text{accept}}$, namely $r<P_{\text{accept}}$, the new solution is also accepted.
@@ 4 Experiments
We assess the capabilities of DCS in reducing class accuracy differences and enhancing overall accuracy with three ICL settings: 1-shot, 5-shot, and $N$-shot. Demonstrations in the prompt are randomly selected from the optimization set for 1-shot and 5-shot cases, while the $N$-shot demonstration cascades examples from each class, to minimize under-representations of any class. The proposed method is compared against SoTA methods, DNIP and FuRud. We prompt LLMs with an A100 GPU to obtain initial ICL outputs, following DNIP’s hyperparameters, then we run simulated annealing on CPUs. Evaluations are performed on seven multi-class text classification tasks across general and biomedical domains, including AGNews (AGN, 4-class;[38]), DBpedia (DBP, 14-class;[1]), SST-5 (SST, 5-class;[39]), TREC (6-class;[40, 41]), RTE (binary;[42]), DDI (5-class;[43]), and PubMedQA (PMQA, 3-class;[44]). Appendix C presents more details.
@@ 4.1 Main Results
The proposed method achieves SoTA overall accuracy improvements on all tasks except for RTE across all ICL settings. As for RTE, due to its simplicity as a binary classification task, all three methods obtain similar accuracy and COBias improvements. Meanwhile, it achieves SoTA COBias performance on general domain tasks. In general, its COBias is on par with the current best methods, with average COBias over all seven tasks falling between that of DNIP and FuRud. On AGNews, DBpeida, and TREC, the proposed method achieves both more overall accuracy improvements and more COBias reduction than DNIP and FuRud across all ICL settings, demonstrating the superior performance gains achieved by ensembling class-level and sample-level debiasing. Weak classes are elevated. We show in Section 4.2 that DCS powers up the low-accuracy classes. We observe that weaker classes are often corrected by rules, while stronger classes are most likely corrected by weights, further highlighting the need of an ensemble method.
@@
Figure 3 plots the time used to optimize correction functions (left axis) and the search space size (right axis) for each task. For fair comparisons, we fix the weight scale for weight correction to 30 alongside 19 triangular membership functions, making a total of 49 possible correction functions and $49\cdot N$ possible assignments across all variables. DCS is computationally efficient where all tasks finish optimization within 150 outer loops, ranging from a maximum of 96.6 seconds to a minimum of 5.4 seconds. Notably, SST-5, DDI, and AGNews have a relatively longer time due to having a larger optimization set (Appendix C). In general, although larger optimization sets and search spaces need more time to converge, the annealing steps needed for a task can be further tuned to achieve cost-effective gains on downstream applications.